6. If we do not know σ=1.20 in the aforementioned problem 4, we will use a T-procedure to perform the hypothesis test. Below no 6
Answer #1

(a)

Sample Mean = \bar{x} = 6.590565

(b)

t score = - 2.3134

ndf =n - 1 = 53 - 1 = 52

Two Tail Test

By Technology, P - Value < 0.0001

Since P - value is less than \alpha = 0.01, the difference is significant.Reject null hypothesis.

Conclusion:
The data support the claim that the population mean is different from 7.

(c)

t score = - 2.3134

ndf =n - 1 = 53 - 1 = 52

One Tail Left Side Test

By Technology, P - Value < 0.0001

Since P - value is less than \alpha = 0.01, the difference is significant.Reject null hypothesis.

Conclusion:
The data support the claim that the population mean is less than - 7.